a new community center is being built in pagosa springs. The perimeter of the rectangular playing field is 312 yards. the lenght of the field is 3 yards less than double the width. what are the dimensions of the playing field?
Answer by Boreal(15235)

width=x
length=2x-3
Half the perimeter is L+W=156
so x+2x-3=156
3x=159
x=53 yards width
2x-3=103 yards length
2W=106
2L=206
it is 103 yards x 53 yards